OpenStudy (vcabral1134):

What is the distance between points (−2, 8) and (6, −2)? Round to the nearest tenth of a unit. units

OpenStudy (vcabral1134):

@confluxepic

OpenStudy (vcabral1134):

@Preetha

OpenStudy (vcabral1134):

Will fan and medal

OpenStudy (vcabral1134):

@iambatman

OpenStudy (vcabral1134):

@sleepyjess

OpenStudy (vcabral1134):

@sarimari

OpenStudy (vcabral1134):

Anybody???

OpenStudy (michele_laino):

we have to apply this formula: \[\Large d = \sqrt {{{\left( {{x_2} - {x_1}} \right)}^2} + {{\left( {{y_2} - {y_1}} \right)}^2}} \]

OpenStudy (michele_laino):

where: (x_1,y_1)=(-2,8), and (x_2,y_2)=(6,-2) d is the requested distance

OpenStudy (vcabral1134):

164?

OpenStudy (michele_laino):

better is: \[d = \sqrt {164} = ...?\]

OpenStudy (vcabral1134):

12.8

OpenStudy (michele_laino):

that's right!

OpenStudy (vcabral1134):

Yay! Thank you for helping me!

OpenStudy (michele_laino):

Thank you! :)